A newspaper on January 24, 1997, discussed the recommendation of a special panel concerning mammograms for women in their 40s. About 2% of women aged 40 to 49 years old develop breast cancer in their 40s. But the mammogram used for women in that age group has a high rate of false positives and false negatives, the false positive rate is 0.22, and the false negative rate is 0.24. If a woman in her 40s has a positive mammogram, what is the probability that she actually has breast cancer? Which of the following formulas will be used to find the probability? Pr(cancer).Pr testlcancer) Pr(cancer). Prtest cancer). Pr(no cancer). Pr(+test no cancer) Pr(cancer). Pr( testcancer) Pr(cancer). Pr+test cancer) Pr(no cancer). Pr(test no cancer) O C. Pr+testjcancer)+ Pr+test no cancer) + Pr-test cancer) + Pr-test no cancer) Pr(cancer).Pr+test cancer) Pr(cancer). Pr+test cancer) Pr(no cancer). Pr(test no cancer) OD. The probability that a woman that has a positive mammogram actually has cancer is Type an integer or a simplified fraction) A newspaper on January 24, 1997, discussed the recommendation of a special panel concerning mammograms for women in their 40s. About 2 % of women aged 40 to 49 years old develop breast cancer in their 40s. But the mammogram used for women in that age group has a high rate of false positives and false negatives; the false positive rate is 0.22 , and the false negative rate is 0.24. If a woman in her 40s has a positive mammogram, what is the probability that she actually has breast cancer?
Answer #1

Ans:

Given that

P(cancer)=0.02

P(no cancer)=1-0.02=0.98

P(+test/no cancer)=0.22

P(-test/cancer)=0.24

Now,

P(+test)=P(+test/cancer)*P(cancer)+P(+test/no cancer)*P(no cancer)

=(1-0.24)*0.02+0.22*0.98

=0.2308

P(cancer/+test)=P(+test/cancer)*P(cancer)/P(+test)

=P(+test/cancer)*P(cancer)/(P(+test/cancer)*P(cancer)+P(+test/no cancer)*P(no cancer))

Option D is correct.

P(cancer/+test)=(1-0.24)*0.02/0.2308=0.0659